\begin{thebibliography}{99}

\bibitem{HESS}
F. Aharonian, A. G. Akhperjanian, K.M. Aye et al.,
\emph{Very high-energy gamma rays from the direction of Sagittarius A*},
\emph{A\&A} 425 (2004) L13;
%F. Aharonian, A. G. Akhperjanian, K.M. Aye et al.
\emph{Spectrum and variability of the Galactic Center VHE gamma-ray source HESS J1745-290},
\emph{A\&A} 503 (2009) 817.

\bibitem{Cembranos:2012nj}
  J.~A.~R.~Cembranos, V.~Gammaldi and A.~L.~Maroto,
  \emph{Possible dark matter origin of the gamma ray emission from the galactic center observed by HESS},
  \emph{Phys.\ Rev.\ D} {\bf 86} (2012) 103506 [arXiv:1204.0655 [hep-ph]]; %%CITATION = ARXIV:1204.0655;%%;
  \emph{Spectral Study of the HESS J1745-290 Gamma-Ray Source as Dark Matter Signal},
  \emph{JCAP} {\bf 1304} (2013) 051  [arXiv:1302.6871 [astro-ph.CO]]. %%CITATION = ARXIV:1302.6871;%%

\bibitem{Belikov:2012ty}
A.~V.~Belikov et al., %, G.~Zaharijas and J.~Silk,
\emph{Study of the Gamma-ray Spectrum from the Galactic Center in view of Multi-TeV Dark Matter Candidates},
\emph{Phys.\ Rev.\ D} {\bf 86} (2012) 083516 [arXiv:1207.2412 [astro-ph.HE]]. %%CITATION = ARXIV:1207.2412;%%

\bibitem{Fermi}
V.~Vitale, A.~Morselli and f.~t.~F.~/L.~Collaboration,
\emph{Indirect Search for Dark Matter from the center of the Milky Way with the Fermi-Large Area Telescope},
arXiv:0912.3828 [astro-ph.HE];
M. Chernyakova {\em et.~al.},
\emph{The high-energy, Arcminute-scale galactic center gamma-ray source},
\emph{ApJ} {\bf 726} (2011) 60.

\bibitem{Ce10}
J.~A.~R.~Cembranos et al., %, A.~de la Cruz-Dombriz, A.~Dobado, R.~Lineros and A.~L.~Maroto,
\emph{Photon spectra from WIMP annihilation},
\emph{Phys.\ Rev.\  D} {\bf 83} (2011) 083507 [arXiv:1009.4936 [hep-ph]];  %%CITATION = PHRVA,D83,083507;%%24
\emph{Photon spectra from quark generation by WIMPs},
\emph{AIP Conf.\ Proc.} {\bf 1343} (2011) 595-597 [arXiv:1011.2137 [hep-ph]]; %%CITATION = ARXIV:1011.2137;%%;
\emph{Fitting formulae for photon spectra from WIMP annihilation},
\emph{J.\ Phys.\ Conf.\ Ser.} {\bf 314} (2011) 012063 [arXiv:1012.4473 [hep-ph]]; %%CITATION = ARXIV:1012.4473;%%;
\emph{Reliability of Monte Carlo event generators for gamma ray dark matter searches},
\emph{JHEP} {\bf 1309} (2013) 077 [arXiv:1305.2124 [hep-ph]]. %%CITATION = ARXIV:1305.2124;%%

\bibitem{Blumenthal}
G.R. Blumenthal, S.M. Faber, R. Flores, J. R. Primack,
\emph{Contraction of Dark Matter Galactic Halos Due to Baryonic Infall},
\emph{ApJ}, {\bf 301} (1986) 27;
F.~Prada et al., %, A.~Klypin, J.~Flix Molina, M.~Mart\'inez, E.~Simonneau,
\emph{Dark Matter Annihilation in the Milky Way Galaxy: Effects of Baryonic Compression}
\emph{Phys.\ Rev.\ Lett.}, {\bf 93} (2004) 241301 [astro-ph/0401512].

\bibitem{branons}
A.~Dobado and A.~L.~Maroto,
\emph{The Dynamics of the Goldstone bosons on the brane},
\emph{Nucl.\ Phys.\ B} {\bf 592} (2001) 203 [hep-ph/0007100]; %%CITATION = HEP-PH/0007100;%%
J.~A.~R.~Cembranos, A.~Dobado and A.~L.~Maroto,
\emph{Brane world dark matter},
\emph{Phys.\ Rev.\ Lett.}  {\bf 90} (2003) 241301 [hep-ph/0302041]; %%CITATION = PRLTA,90,241301;%%
\emph{Cosmological and astrophysical limits on brane fluctuations},
\emph{Phys.\ Rev. D} {\bf 68} (2003) 103505 [hep-ph/0307062]; %%CITATION = PHRVA,D68,103505;%%cos
\emph{Dark geometry},
\emph{Int. J. Mod. Phys.} {\bf D13} (2004) 2275 [hep-ph/0405165]; %%CITATION = IMPAE,D13,2275;%%DG
A.~L.~Maroto,
\emph{The Nature of branon dark matter},
\emph{Phys.\ Rev.\ D} {\bf 69} (2004) 043509 [hep-ph/0310272]; %%CITATION = PHRVA,D69,043509;%%
\emph{Brane oscillations and the cosmic coincidence problem},
\emph{Phys.\ Rev.\ D} {\bf 69} (2004) 101304 [hep-ph/0402278]; %%CITATION = PHRVA,D69,101304;%%
J.~A.~R.~Cembranos  {\it et al.},
\emph{Is the CMB Cold Spot a gate to extra dimensions?},
\emph{JCAP} {\bf 0810} (2008) 039 [0803.0694 [astro-ph]]. %%CITATION = JCAPA,0810,039;%%17

\bibitem{branonsgamma}
J. A. R. Cembranos, A. de la Cruz-Dombriz, V. Gammaldi, A.L. Maroto,
\emph{Detection of branon dark matter with gamma ray telescopes},
\emph{Phys. Rev. D} 85 (2012) 043505 arXiv:1111.4448 [astro-ph.CO]. %%CITATION = ARXIV:1111.4448;%%

\bibitem{lab}
J.~Alcaraz {\it et al.},
\emph{Limits on the brane fluctuations mass and on the brane tension scale from electron positron colliders},
\emph{Phys. Rev. D} {\bf 67} (2003) 075010 [hep-ph/0212269]; %%CITATION = PHRVA,D67,075010;%%
P. Achard {\it et al.},
\emph{Search for branons at LEP},
\emph{Phys. Lett.} {\bf B597} (2004) 145 [hep-ex/0407017]; %%CITATION = PHLTA,B597,145;%%
J.~A.~R.~Cembranos, A.~Dobado and A.~L.~Maroto,
\emph{Brane skyrmions and wrapped states},
\emph{Phys. Rev. D} {\bf 65} (2002) 026005 [hep-ph/0106322]; %%CITATION = PHRVA,D65,026005;%%
\emph{Branon search in hadronic colliders},
\emph{Phys. Rev. D} {\bf 70} (2004) 096001 [hep-ph/0405286]; %%CITATION = PHRVA,D70,096001;%%22
\emph{Branon radiative corrections to collider physics and precision observables},
\emph{Phys.\ Rev.\ D} {\bf 73} (2006) 035008 [hep-ph/0510399]; %%CITATION = PHRVA,D73,035008;%%rad
\emph{Dark matter clues in the muon anomalous magnetic moment},
\emph{Phys.\ Rev.\ D} {\bf 73} (2006) 057303 [hep-ph/0507066]; %%CITATION = PHRVA,D73,057303;%%rad
\emph{Some model-independent phenomenological consequences of flexible brane worlds},
\emph{J.\ Phys.\ A}  {\bf 40} (2007) 6631 [hep-ph/0611024]; %%CITATION = JPAGB,A40,6631;%%
J.~A.~R.~Cembranos, J.~L.~Diaz-Cruz and L.~Prado,
\emph{Impact of DM direct searches and the LHC analyses on branon phenomenology},
\emph{Phys.\ Rev.\ D} {\bf 84} (2011) 083522 [arXiv:1110.0542 [hep-ph]]; %%CITATION = ARXIV:1110.0542;%%
J.~A.~R.~Cembranos, R.~L.~Delgado and A.~Dobado,
\emph{Brane-Worlds at the LHC: Branons and KK-gravitons},
arXiv:1306.4900 [hep-ph]. %%CITATION = ARXIV:1306.4900;%%

\bibitem{cosmics}
S.~Rudaz and F.~W.~Stecker,
\emph{Cosmic ray anti-protons, positrons and gamma-rays from halo dark matter annihilation},
\emph{Astrophys.\ J.}  {\bf 325} (1988) 16;
J.~A.~R.~Cembranos, J.~L.~Feng, A.~Rajaraman and F.~Takayama,
\emph{SuperWIMP solutions to small scale structure problems},
\emph{Phys.\ Rev.\ Lett.}  {\bf 95} (2005) 181301 [hep-ph/0507150]; %%CITATION = PRLTA,95,181301;%%1
J.~A.~R.~Cembranos, J.~L.~Feng and L.~E.~Strigari,
\emph{Resolving Cosmic Gamma Ray Anomalies with Dark Matter Decaying Now},
\emph{Phys.\ Rev.\ Lett.}  {\bf 99} (2007) 191301 [arXiv:0704.1658 [astro-ph]]; %%CITATION = PRLTA,99,191301;%%32
\emph{Exotic Collider Signals from the Complete Phase Diagram of Minimal Universal Extra Dimensions},
\emph{Phys.\ Rev.\  D} {\bf 75} (2007) 036004 [hep-ph/0612157]; %%CITATION = PHRVA,D75,036004;%%2
J.~A.~R.~Cembranos and L.~E.~Strigari,
\emph{Diffuse MeV Gamma-rays and Galactic 511 keV Line from Decaying WIMP Dark Matter},
\emph{Phys.\ Rev.\  D} {\bf 77} (2008) 123519 [arXiv:0801.0630 [astro-ph]]; %%CITATION = PHRVA,D77,123519;%%23
J.~A.~R.~Cembranos,
\emph{The Newtonian limit at intermediate energies},
\emph{Phys.\ Rev.\  D} {\bf 73} (2006) 064029 [arXiv:gr-qc/0507039]; %%CITATION = PHRVA,D73,064029;%%6
\emph{Dark Matter from R2-gravity},
\emph{Phys.\ Rev.\ Lett.}  {\bf 102} (2009) 141301 [arXiv:0809.1653 [hep-ph]]. %%CITATION = PRLTA,102,141301;%%5

\end{thebibliography}

\documentclass{PoS}

\title{Cosmic Rays from Heavy Dark Matter from the Galactic Center}

\ShortTitle{Cosmic Rays from Heavy Dark Matter from the Galactic Center}

\author{\speaker{Jose A. R. CEMBRANOS}
         \\
        Departamento de  F\'{\i}sica Te\'orica I, Universidad Complutense de Madrid, E-28040 Madrid, Spain\\
        E-mail: \email{cembra@fis.ucm.es}}

\author{Viviana GAMMALDI\\
        Departamento de  F\'{\i}sica Te\'orica I, Universidad Complutense de Madrid, E-28040 Madrid, Spain\\
        E-mail: \email{vivigamm@pas.ucm.es}}

\author{Antonio L. MAROTO\\
        Departamento de  F\'{\i}sica Te\'orica I, Universidad Complutense de Madrid, E-28040 Madrid, Spain\\
        E-mail: \email{maroto@fis.ucm.es}}

\abstract{
The gamma-ray fluxes observed by the High Energy Stereoscopic System (HESS) from the J1745-290 Galactic Center source is well fitted by the secondary photons coming from Dark Matter (DM) annihilation in particle-antiparticle standard model pairs over a diffuse power-law background. The spectral features of the signal are consistent with different channels: light quarks, electro-weak gauge bosons and top-antitop production. The amount of photons and morphology of the signal localized within a region of few parsecs, require compressed DM profiles as those resulting from baryonic contraction, which offer large enhancements in the signal over DM alone simulations. The fits return a heavy WIMP, with a mass above 10 TeV, but well below the unitarity limit for thermal relic annihilation.
The fitted background spectral index is compatible with the Fermi-Large Area Telescope (LAT) data from the same region. This possibility can be potentially tested with the observations of other high energy cosmic rays.
}

\FullConference{The European Physical Society Conference on High Energy Physics \\
                 18-24 July, 2013\\
                 Stockholm, Sweden}

\begin{document}

\section{Gamma rays from the galactic center and the dark matter hypothesis}

We have studied the possibility of explaining the gamma ray data \cite{HESS} observed by the High Energy Stereoscopic System (HESS) from the central part of
our galaxy by being partially produced by Dark Matter (DM) annihilations or decays \cite{Cembranos:2012nj,Belikov:2012ty}.
The complexity of the region and the amount of
delocalized emitting sources justifies the hypothesis of a non-negligible background. DM annihilations or decays into single standard model
particle-antiparticle channels provide good fits if the DM signal is complemented with such a background, that it is consistent with Fermi-LAT
measurements \cite{Fermi}.

The fits return a DM mass between $15$ and $110 \; \rm{TeV}$ depending on the studied channel \cite{Ce10}. Leptonic channels
are clearly disfavored, but hadronic channels such us the $d\bar d$, or electroweak channels such as the $W^+W^-$ and $ZZ$ channels,
offer very good consistency with $\chi^2/dof=0.73$, $0.84$ and $0.85$ respectively \cite{Cembranos:2012nj}.
The morphology of the signal requires compressed dark halos
as the ones that take into account baryonic dissipation \cite{Blumenthal}.

A motivated DM candidate which could have high enough masses and account for the right abundance of DM as a thermal relic,
is the branon field \cite{branons,branonsgamma}, which corresponds to brane fluctuations in flexible brane-world models.
We have proved that a branon mass of $M \simeq 50.6$ TeV provides a good fit to the mentioned HESS data \cite{Cembranos:2012nj}.
%At these masses, its main contribution to gamma rays comes from the annihilation into gauge bosons.

\section{Conclusions and future work}

The heavy DM masses required for fitting the HESS data are practically unconstrained by direct detection searches
or colliders experiments \cite{lab}.
On the contrary, this possibility can be tested with the observations of other cosmic-rays \cite{cosmics} from the GC and from other astrophysical objects.
For example, high energy neutrinos could confirm this hypothesis at IceCube or ANTARES for angular resolutions below one degree.

\acknowledgments{
This work has been supported by MICINN (Spain) project numbers FIS2011-23000, FPA2011-27853-01 and Consolider-Ingenio MULTIDARK CSD2009-00064.
}

\end{document}